How do you prepare 500 mL of 250 mM Tris base? (The molecular weight of Tris base is 121.14 g/mol)
Answer #1

Molarity = Moles/volume of solution in liter

  • Given molarity = 250 mM = 250/1000 M = 0.25 M
  • Volume = 500 ml = 0.5 L
  • Moles = 0.25 × 0.5 = 0.125

Moles = given mass/ molar mass

Mass of tried base required = 0.125 × 121.14 = 15.1425 g

Therefore to prepare 250 mM of 500 ml of Tried base solution, we mix 15.425 g of tried base in 500 ml of solution

Feel free to comment if you have any problem in understanding this answer.